OpenStudy (anonymous):

the sum of two numbers re 145. the smaller number is decreased by 2 and the larger number is increased by 1. their ratio are 2:7. Find the two numbers ??? .. can you plss help me solving itt?? :((

OpenStudy (anonymous):

pls help me :((

OpenStudy (jiteshmeghwal9):

Let the smaller number be x. then Larger number=145-x. A/q, smaller number=x-2 Larger number=145-x+1 =146-x\[\Large{\frac{x-2}{146-x}=\frac{2}{7}}\] solve for 'x' by cross multiplying :)

terenzreignz (terenzreignz):

@jiteshmeghwal9 I think it should be x for the smaller number and 145 - x for the larger number, as you originally stated.

OpenStudy (anonymous):

x= 296.56

OpenStudy (jiteshmeghwal9):

yeah sorry :)

OpenStudy (anonymous):

so, how to solve very clearly?

OpenStudy (jiteshmeghwal9):

u want full solution step by step ???

OpenStudy (jiteshmeghwal9):

@Jezrielanides

OpenStudy (anonymous):

sure :)

OpenStudy (anonymous):

very NEEDED :)) thank you very much Friends :)

OpenStudy (jiteshmeghwal9):

Let the smaller number be x. then Larger number=145-x. A/q, smaller number=x-2 Larger number=145-x+1 =146-x \[\Large{\frac{x-2}{146-x}=\frac{2}{7}}\]by cross multiplying\[\Large{7(x-2)=2(146-x)}\]solve this for 'x' & then put the value in the equations (x) for smaller number & (145-x) for larger number :)

OpenStudy (anonymous):

@jiteshmeghwal9 the X is 34 or x=34 .. then?

OpenStudy (anonymous):

AAAAHAHH!! so it is enough now!! thank you I understand it clearly !! :)))))))))))))))))))))